Consider the following bond: Face value = 1000; coupon rate = 8%; maturity = 5 years; ytm = 7%

A) What is the value of the bond today and in 2 years?

b) what are the current yield and capital gains yield for this bond this year and in two years?

c) Assuming interest rates remain the same over this bond's lifetime, what is going to happen to the value of this bond as time goes by?
Answer #1

a) Face value = 1000
maturity = 5 years; ytm = 7%
Coupon =8%*1000 =80
Price of bond today =PV of Coupons+PV of Par Value =80*((1-(1+7%)^-5)/7%+1000/(1+7%)^5 =1041.02

Price of bond in 2 years =PV of Coupons+PV of Par Value =80*((1-(1+7%)^-3)/7%+1000/(1+7%)^3 =1026.24

Price of bond in 1 year =PV of Coupons+PV of Par Value =80*((1-(1+7%)^-4)/7%+1000/(1+7%)^4 =1033.87

b)Current Yield =Coupon/Price =80/1041.02=7.68%
Capital Gain Yield =(1033.87-1041.02)/1041.02=-0.68%

Current Yield in 2 years =80/1026.24 =7.80%
Capital gain yield =YTM -Current Yield =7%-7.80% =-0.80%

c) As lifetime increases value of bond decreases.